For the following reaction, 4.73 grams of chlorine gas are mixed with excess bromine . The reaction yields 14.3 grams of bromine monochloride .
What is the theoretical yield of bromine monochloride ?
What is the percent yield ?
1
Expert's answer
2017-11-22T14:03:07-0500
Reaction of bromine monochloride’s formation:
〖Br〗_2+〖Cl〗_2=2BrCl
Theoretical mass of bromine monochloride can be calculated by the following manner:
n(〖Cl〗_2 )=2n(BrCl)
m(〖Cl〗_2 )/M(〖Cl〗_2 ) =2m(BrCl)/M(BrCl)
m(BrCl)=(m(〖Cl〗_2 )∙M(BrCl))/(2∙M(〖Cl〗_2 ) )=(4.73g ∙115.4)/(2∙35.5)=7.69 g
The percent yield:
η=m_pract/m_theor 100%=14.3/7.69 100%=194%
Normally, percent yield should be less than 100%. In obtained case the percent yield is 194%, that is why the possibility of incorrect data from the question take place.
Answer:
Theoretical mass of bromine monochloride is 7.69 g, percent yield is 194%.